# The Quantum Decoding Problem : Tight Achievability Bounds and Application to Regev’s Reduction

### v1

@@ 3.2 Achievability Result
Fix $R\in(0,H)$ and let $k_{i}=\lfloor Rn_{i}\rfloor$. From Proposition 5 we know that the average probability $\overline{\textrm{P}_{\textrm{PGM}}}$ over the randomness of codes (obtained by choosing a generator matrix $\mathbf{G}$ of $\mathscr{C}$} at random) is given by
$\overline{\textrm{P}_{\textrm{PGM}}}$ $=$ $\frac{1}{q^{k_{i}}}\mathbb{E}_{\mathbf{G}}\left(\sum\limits_{{\boldsymbol{s}}\in\mathbb{F}_{q}^{n}/{\mathscr{C}}^{\perp}}Z_{\boldsymbol{s}}\right)^{2}\quad\text{where}$ (8)
@@
When either ${\boldsymbol{x}}-{\boldsymbol{s}}$ or ${\boldsymbol{y}}-{\boldsymbol{s}}$ is zero:
@@
In this case $f_{i}=g^{\otimes i}$ and we take $n_{i}=i$, $H=H_{q}(|\widehat{g}|^{2})$, $A_{i}(\epsilon)=q^{-i(H+\epsilon)}$, $B_{i}(\epsilon)=q^{-i(H-\epsilon)}$. Proposition 2 shows that $\delta(\epsilon,n)=\operatorname{negl}(n)$ for all $\epsilon>0$. This shows that we can apply Theorem 4 to this case, from which we derive immediately Theorem 1. ∎
@@ 3.3 Strong converse
The proof of this statement will rely on several statements. The first one uses the fact that for any $\boldsymbol{c}$} in $\mathscr{C}_{i}$ the state $\widehat{|\psi_{{\boldsymbol{c}}}\rangle}$ which is equal to $\sum_{{\boldsymbol{y}}\in\mathbb{F}_{q}^{n_{i}}}\widehat{f}_{i}({\boldsymbol{y}})\chi_{{\boldsymbol{c}}}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle$ by Lemma 3 is extremely close to the state $|\widetilde{\psi}_{\boldsymbol{c}}\rangle$ which is obtained from $\widehat{|\psi_{{\boldsymbol{c}}}\rangle}$ by keeping in the previous sum only those corresponding to a $\boldsymbol{y}$} that are in the typical set ${\mathcal{T}}_{\epsilon(i)}^{(n_{i})}$} associated to $|\widehat{f}_{i}|^{2}$.
@@
For the last point we used Lemma 1 which says that $|{{\mathcal{T}}_{\epsilon(i)}^{(n_{i})}}|\leq\frac{1}{A_{i}(\epsilon(i))}$ and the fact that $A_{i}(\epsilon(i))=q^{-n_{i}(H+o(1))}$. Now, consider a quantum POVM $\{M_{\boldsymbol{c}}\}_{{\boldsymbol{c}}\in{\mathscr{C}}_{i}}$. Let $p_{\textrm{succ}}$} be the average probability to solve the $\mathrm{QDP}$ problem using this measurement whereas $\widetilde{p}$ is the average probability of distinguishing $\boldsymbol{c}$} with the same POVM when fed instead of $|\widehat{\psi}_{\boldsymbol{c}}\rangle$ the state $|\widetilde{\psi}_{\boldsymbol{c}}\rangle$. By using Lemma 10 we deduce that
@@ 3.4 Application to the rank metric
The function $f_{t}$ is radial, meaning that the value $f_{t}(\boldsymbol{e})$ only depends on the rank weight $|\boldsymbol{e}|_{\text{rk}}$ of $\boldsymbol{e}$. We will be interested in the dual distribution $p\mathop{=}\limits^{\triangle}\left|\widehat{f}_{t}\right|^{2}=\left|f_{b-t}\right|^{2}$. This probability distribution concentrates around the rank weight ntn-t as shown in Appendix B.2 of [DRT24]
@@ 4 Sampling Dual Codewords
  • at the limit of the tractability regime, we essentially sample the most likely non zero dual codewords according to the aforementioned a probability distribution.
@@ 4.1 The sampling algorithm
In the whole section $\mathscr{C}$ denotes a linear code of length $n$ over $\mathbb{F}_{q}$} whose generator matrix $\mathbf{G}$ is drawn uniformly at random in $\mathbb{F}_{q}^{k\times n}$, meaning that the code is of dimension $k$ with probability 1o(1)1-o(1) for a fixed rate $R=\frac{k}{n}$ bounded away from $1$. Roughly speaking, in order to avoid destroying the second register carrying $|\psi_{\boldsymbol{c}}\rangle$ and still measuring most of the time $\boldsymbol{c}$}, we apply the PGM coherently on the second register carrying $|{\boldsymbol{c}}\rangle$ and write the output on the third register. The value of the third register is then subtracted to the first register and we reverse at that point the PGM between the second and third register. As we have seen in the previous section it is more convenient to express the PGM in this case in the Fourier basis. We apply therefore prior to applying the PGM the quantum Fourier transform on the second register carrying $|\psi_{\boldsymbol{c}}\rangle$. This is basically what is done in Algorithm of Section 6.3 in [CT23], which we recall now Algorithm 4.1 : Algorithm of the quantum reduction with PGM. Initial state preparation: $\quad|\phi_{0}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\sum_{\boldsymbol{e}\in\mathbb{F}_{q}^{n}}f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|\boldsymbol{e}\rangle|{\boldsymbol{0}}\rangle$ adding $\boldsymbol{c}$} to $\boldsymbol{e}$: $\mapsto$ $\quad|\phi_{1}\rangle$ $=$ $\quad\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\sum_{\boldsymbol{e}\in\mathbb{F}_{q}^{n}}f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|{\boldsymbol{c}}+\boldsymbol{e}\rangle|{\boldsymbol{0}}\rangle$ $=$ $\quad\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}|{\boldsymbol{c}}\rangle|\psi_{\boldsymbol{c}}\rangle|{\boldsymbol{0}}\rangle$ apply the QFT on $|\psi_{\boldsymbol{c}}\rangle$: $\mapsto$ $\quad|\phi_{2}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}|{\boldsymbol{c}}\rangle|\widehat{\psi_{\boldsymbol{c}}}\rangle|{\boldsymbol{0}}\rangle$ applying the PGM coherently: $\mapsto$ $\quad|\phi_{3}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}|{\boldsymbol{c}}\rangle\sum_{{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|Y_{{\boldsymbol{c}}^{\prime}}\rangle|{\boldsymbol{c}}^{\prime}\rangle$ $\text{ with }|\widehat{\psi_{\boldsymbol{c}}}\rangle=\sum_{{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|Y_{{\boldsymbol{c}}^{\prime}}\rangle$ erase $\boldsymbol{c}$}: $\mapsto$ $\quad|\phi_{4}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|{\boldsymbol{c}}-{\boldsymbol{c}}^{\prime}\rangle|Y_{{\boldsymbol{c}}^{\prime}}\rangle|{\boldsymbol{c}}^{\prime}\rangle$ reverse PGM: $\mapsto$ $\quad|\phi_{5}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|{\boldsymbol{c}}-{\boldsymbol{c}}^{\prime}\rangle|Y_{{\boldsymbol{c}}^{\prime}}\rangle|{\boldsymbol{0}}\rangle$ $=$ $\quad\sqrt{\textrm{P}_{\textrm{PGM}}}|{\boldsymbol{0}}\rangle\left(\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}|Y_{\boldsymbol{c}}\rangle\right)+\sum_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}\neq{\boldsymbol{c}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|{\boldsymbol{c}}-{\boldsymbol{c}}^{\prime}\rangle|Y_{{\boldsymbol{c}}^{\prime}}\rangle$ measure 1st register, if $|{\boldsymbol{0}}\rangle$: $\mapsto$ $\quad|\phi_{6}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}|Y_{{\boldsymbol{c}}}\rangle=\dfrac{1}{n_{0}}\sum_{{\boldsymbol{y}}\in{\mathscr{C}}^{\perp}}\hat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle\;\;\text{w.p. $\textrm{P}_{\textrm{PGM}}$}$ measuring the whole state: $\mapsto$ $\quad|{\boldsymbol{y}}\rangle\;\;\text{with ${\boldsymbol{y}}\in{\mathscr{C}}^{\perp}$.}$
@@
4.2 A slight tweak in the algorithm which avoids measuring the zero codeword
The idea is to tweak the algorithm in order to get as the last state before the final measurement the state $|U_{0}\rangle$ rather than $|\phi_{6}\rangle=\dfrac{1}{n_{0}}\sum_{{\boldsymbol{y}}\in{\mathscr{C}}^{\perp}}\hat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle$, where
@@ 4.2 A slight tweak in the algorithm which avoids measuring the zero codeword
$|{Z}_{\boldsymbol{c}}\rangle=\frac{1}{\sqrt{q^{k}}}\left(|U_{0}\rangle+\sum_{{\boldsymbol{s}}\neq\mathbf{0}}\chi_{\boldsymbol{c}}(u_{\boldsymbol{s}})|\widetilde{W}_{\boldsymbol{s}}\rangle\right).$
These states are readily seen to be orthogonal. However projecting in this basis does not make the measurement complete. This can be circumvented by adding to the basis the $|{\boldsymbol{0}}\rangle$ state. We add an extra outcome $\boldsymbol{u}\in\mathbb{F}_{q}^{n}\backslash{\mathscr{C}}$, define ${\mathscr{S}}={\mathscr{C}}\cup\{\boldsymbol{u}\}$ and set $|Z_{\boldsymbol{u}}\rangle=|{\boldsymbol{0}}\rangle$. We therefore have a projective measurement $\{|Z_{{\boldsymbol{y}}}\rangle\}_{{\boldsymbol{y}}\in{\mathscr{S}}}$. Again, we have $\langle\widehat{\psi_{\boldsymbol{c}}}|{Z}_{\boldsymbol{c}}\rangle\geq\sqrt{\textrm{P}_{\textrm{PGM}}}-\frac{1}{\sqrt{q^{k}}}$ and independent of $\boldsymbol{c}$} so w.p. at least $\left(\sqrt{\textrm{P}_{\textrm{PGM}}}-\frac{1}{\sqrt{q^{k}}}\right)^{2}$, we get the state $|U_{0}\rangle$. Then, if we measure this state $|U_{0}\rangle$, we will get a dual codeword $\boldsymbol{y}$} with probability $q({\boldsymbol{y}})$ given by
@@
Let $q\geq 2$ be a prime power. We assume that we have an increasing sequence of positive integers $(n_{i})_{i\geq 1}$ and a family of functions $(f_{i})_{i\geq 1}$ with $f_{i}:\mathbb{F}_{q}^{n_{i}}\rightarrow\mathbb{C}$ with $\lvert\kern-1.0pt\lvert f_{i}\rvert\kern-1.0pt\rvert_{2}=1$ such that the sequence $(p_{i})_{i\geq 1}$ defined by $p_{i}\mathop{=}\limits^{\triangle}|\widehat{f_{i}}|^{2}$ admits a nice family of typical sets not containing 0. We let $H\mathop{=}\limits^{\triangle}\lim_{i\rightarrow\infty}\frac{H_{q}(p_{i})}{n_{i}}$. Let $R$ be a real number chosen in $(0,H)$. Consider a family $({\mathscr{C}}_{i})_{i\geq 1}$ of random linear codes where ${\mathscr{C}}_{i}$ is chosen with a generator matrix chosen uniformly at random in $\mathbb{F}_{q}^{\lfloor Rn_{i}\rfloor\times n_{i}}$. Then, for any $\epsilon>0$ the algorithm of the quantum reduction with PGM using the measurement basis $\{|Z_{{\boldsymbol{y}}}\rangle\}_{{\boldsymbol{y}}\in{\mathscr{S}}}$ described above, outputs codewords of $\mathscr{C}_{i}^{\perp}$ in the typical set ${\mathcal{T}}_{\epsilon}^{(n_{i})}$} with probability $1-\operatorname{negl}(i)$ as $i$ tends to infinity.
@@
Let $g:{\mathbb{F}_{q}}\rightarrow\mathbb{C}$ with $\lvert\kern-1.0pt\lvert g\rvert\kern-1.0pt\rvert_{2}=1$ and $f:\mathbb{F}_{q}^{n}\rightarrow\mathbb{C}$ such that $f=g^{\otimes n}$ where $n$ is an arbitrary positive integer. Let $R$ be a real number in $(0,H_{q}(|\widehat{g}|^{2}))$. Choose a linear code $\mathscr{C}$ at random by picking its generator matrix uniformly at random in $\mathbb{F}_{q}^{\lfloor Rn\rfloor\times n}$. Then, for any $\epsilon>0$ the algorithm of the quantum reduction with PGM using the measurement basis $\{|Z_{{\boldsymbol{y}}}\rangle\}_{{\boldsymbol{y}}\in{\mathscr{S}}}$ described above, outputs codewords of $\mathscr{C}^{\perp}$ in the typical set ${\mathcal{T}}_{\epsilon}^{(n)}$} with probability $1-\operatorname{negl}(n)$ as $n$ tends to infinity.

### v2

@@ 3.2 Achievability Result
Fix $R\in(0,H)$ and let $k_{i}=\lfloor Rn_{i}\rfloor$. From Proposition 5, we know that the average probability $\overline{\textrm{P}_{\textrm{PGM}}}$ over the randomness of codes (obtained by choosing a generator matrix $\mathbf{G}$ of $\mathscr{C}$} at random) is given by
$\overline{\textrm{P}_{\textrm{PGM}}}$ $=$ $\frac{1}{q^{k_{i}}}\mathbb{E}_{\mathbf{G}}\left(\sum\limits_{{\boldsymbol{s}}\in\mathbb{F}_{q}^{n_{i}}/{\mathscr{C}}^{\perp}}Z_{\boldsymbol{s}}\right)^{2}\quad\text{where}$ (11)
@@
When either ${\boldsymbol{x}}={\boldsymbol{s}}$ or ${\boldsymbol{y}}={\boldsymbol{s}}$:
@@
In this case, $f_{i}=g^{\otimes i}$ and we take $n_{i}=i$, $H=H_{q}(|\widehat{g}|^{2})$, $A_{i}(\epsilon)=q^{-i(H+\epsilon)}$, $B_{i}(\epsilon)=q^{-i(H-\epsilon)}$. Proposition 2 shows that $\delta(\epsilon,n)=\operatorname{negl}(n)$ for all $\epsilon>0$. This shows that we can apply Theorem 5 to this case, from which we derive immediately Theorem 1. ∎
@@ 3.3 Strong converse
The proof of this statement relies on several statements. The first one uses the fact that for any $\boldsymbol{c}$} in $\mathscr{C}_{i}$ the state $\widehat{|\psi_{{\boldsymbol{c}}}\rangle}$ which is equal to $\sum_{{\boldsymbol{y}}\in\mathbb{F}_{q}^{n_{i}}}\widehat{f}_{i}({\boldsymbol{y}})\chi_{{\boldsymbol{c}}}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle$ by Lemma 3 is extremely close to the state $|\widetilde{\psi}_{\boldsymbol{c}}\rangle$ which is obtained from $\widehat{|\psi_{{\boldsymbol{c}}}\rangle}$ by keeping in the previous sum only those corresponding to a $\boldsymbol{y}$} that are in the typical set ${\mathcal{T}}_{\epsilon(i)}^{(n_{i})}$} associated to $|\widehat{f}_{i}|^{2}$.
@@
For the last point, we used Lemma 1 which says that $|{{\mathcal{T}}_{\epsilon(i)}^{(n_{i})}}|\leq\frac{1}{A_{i}(\epsilon(i))}$ and the fact that $A_{i}(\epsilon(i))=q^{-n_{i}(H+o(1))}$. Now, consider a quantum POVM $\{M_{\boldsymbol{c}}\}_{{\boldsymbol{c}}\in{\mathscr{C}}_{i}}$. Let $p_{\textrm{succ}}$} be the average probability to solve the $\mathrm{QDP}$ problem using this measurement whereas $\widetilde{p}$ is the average probability of distinguishing $\boldsymbol{c}$} with the same POVM when fed instead of $|\widehat{\psi}_{\boldsymbol{c}}\rangle$ the state $|\widetilde{\psi}_{\boldsymbol{c}}\rangle$. By using Lemma 10 we deduce that
@@ 3.4 Application to the rank metric
The function $f_{t}$ is radial, meaning that the value $f_{t}(\boldsymbol{e})$ only depends on the rank weight $|\boldsymbol{e}|_{\text{rk}}$ of $\boldsymbol{e}$. We will be interested in the dual distribution $p\mathop{=}\limits^{\triangle}\left|\widehat{f}_{t}\right|^{2}=\left|f_{b-t}\right|^{2}$. This probability distribution concentrates around the rank weight btb-t as shown in Appendix B.2 of [DRT24]
@@ 4 Sampling Dual Codewords
  • at the limit of the tractability regime, we essentially sample the most likely non zero dual codewords according to the aforementioned probability distribution.
@@ 4.1 The sampling algorithm
In the whole section $\mathscr{C}$ denotes a linear code of length $n$ over $\mathbb{F}_{q}$} whose generator matrix $\mathbf{G}$ is drawn uniformly at random in $\mathbb{F}_{q}^{k\times n}$, meaning that the code is of dimension $k$ with probability 1o(1)1-o(1) for a fixed rate $R=\frac{k}{n}$ bounded away from $1$. Roughly speaking, in order to avoid destroying the second register carrying $|\psi_{\boldsymbol{c}}\rangle$ and still measuring most of the time $\boldsymbol{c}$}, we apply the PGM coherently on the second register carrying $|{\boldsymbol{c}}\rangle$ and write the output on the third register. In our case applying the PGM coherently basically amounts to apply the following transformation
$|\widehat{\psi_{\boldsymbol{c}}}\rangle|{\boldsymbol{0}}\rangle\stackrel{{\scriptstyle\text{PGM}}}{{\mapsto}}\sum_{{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|Y_{{\boldsymbol{c}}^{\prime}}\rangle|{\boldsymbol{c}}^{\prime}\rangle$ (30)
where $|\widehat{\psi_{\boldsymbol{c}}}\rangle=\sum_{{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|Y_{{\boldsymbol{c}}^{\prime}}\rangle$ and the $|Y_{\boldsymbol{c}}\rangle$ for $\boldsymbol{c}$} in $\mathscr{C}$} form an orthonormal set of vectors defined by
$|Y_{\boldsymbol{c}}\rangle\mathop{=}\limits^{\triangle}\frac{1}{\sqrt{q^{k}}}\sum_{{\boldsymbol{s}}\in\mathbb{F}_{q}^{k}}\chi_{\boldsymbol{c}}(\boldsymbol{u}_{\boldsymbol{s}})|\tilde{W_{\boldsymbol{s}}}\rangle$ (31)
where $\boldsymbol{u}_{\boldsymbol{s}}$ is an arbitrary element of ${\mathscr{C}}_{\boldsymbol{s}}^{\perp}\mathop{=}\limits^{\triangle}\{{\boldsymbol{x}}\in\mathbb{F}_{q}^{n}:\mathbf{G}{{\boldsymbol{x}}}^{\intercal}={{\boldsymbol{s}}}^{\intercal}\}$ and
$|\tilde{W}_{\boldsymbol{s}}\rangle$ $\mathop{=}\limits^{\triangle}$ $\frac{|W_{\boldsymbol{s}}\rangle}{Z_{\boldsymbol{s}}}\text{ with}$ (32)
$|W_{\boldsymbol{s}}\rangle$ $\mathop{=}\limits^{\triangle}$ $\sum_{{\boldsymbol{y}}\in{\mathscr{C}}_{\boldsymbol{s}}^{\perp}}\widehat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle$ (33)
$Z_{\boldsymbol{s}}$ $\mathop{=}\limits^{\triangle}$ $\lvert\kern-1.0pt\lvert Z_{\boldsymbol{s}}\rvert\kern-1.0pt\rvert$. (34)
It turns out, see Proposition 18 in [CT23] that the PGM associated to the ensemble of states $\{\widehat{|\psi_{\boldsymbol{c}}\rangle}\}_{{\boldsymbol{c}}\in{\mathscr{C}}}$ is nothing but the projective measurement $\{|Y_{\boldsymbol{c}}\rangle\langle Y_{\boldsymbol{c}}|\}_{{\boldsymbol{c}}\in{\mathscr{C}}}$. This explains (30). The value of the third register is then subtracted to the first register and we reverse at that point the PGM between the second and third register. As we have seen in the previous section it is more convenient to express the PGM in this case in the Fourier basis. We apply therefore prior to applying the PGM the quantum Fourier transform on the second register carrying $|\psi_{\boldsymbol{c}}\rangle$. This is basically what is done in Algorithm of Section 6.3 in [CT23], which we now recall Algorithm 4.1 : Algorithm of the quantum reduction with PGM. Initial state preparation: $\quad|\phi_{0}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\sum_{\boldsymbol{e}\in\mathbb{F}_{q}^{n}}f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|\boldsymbol{e}\rangle|{\boldsymbol{0}}\rangle$ adding $\boldsymbol{c}$} to $\boldsymbol{e}$: $\mapsto$ $\quad|\phi_{1}\rangle$ $=$ $\quad\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}\sum_{\boldsymbol{e}\in\mathbb{F}_{q}^{n}}f(\boldsymbol{e})|{\boldsymbol{c}}\rangle|{\boldsymbol{c}}+\boldsymbol{e}\rangle|{\boldsymbol{0}}\rangle$ $=$ $\quad\frac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}|{\boldsymbol{c}}\rangle|\psi_{\boldsymbol{c}}\rangle|{\boldsymbol{0}}\rangle$ apply the QFT on $|\psi_{\boldsymbol{c}}\rangle$: $\mapsto$ $\quad|\phi_{2}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}|{\boldsymbol{c}}\rangle|\widehat{\psi_{\boldsymbol{c}}}\rangle|{\boldsymbol{0}}\rangle$ applying the PGM coherently: $\mapsto$ $\quad|\phi_{3}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}|{\boldsymbol{c}}\rangle\sum_{{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|Y_{{\boldsymbol{c}}^{\prime}}\rangle|{\boldsymbol{c}}^{\prime}\rangle$ $\text{ with }|\widehat{\psi_{\boldsymbol{c}}}\rangle=\sum_{{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|Y_{{\boldsymbol{c}}^{\prime}}\rangle$ erase $\boldsymbol{c}$}: $\mapsto$ $\quad|\phi_{4}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|{\boldsymbol{c}}-{\boldsymbol{c}}^{\prime}\rangle|Y_{{\boldsymbol{c}}^{\prime}}\rangle|{\boldsymbol{c}}^{\prime}\rangle$ reverse PGM: $\mapsto$ $\quad|\phi_{5}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}\in{\mathscr{C}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|{\boldsymbol{c}}-{\boldsymbol{c}}^{\prime}\rangle|Y_{{\boldsymbol{c}}^{\prime}}\rangle|{\boldsymbol{0}}\rangle$ $=$ $\quad\sqrt{\textrm{P}_{\textrm{PGM}}}|{\boldsymbol{0}}\rangle\left(\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}|Y_{\boldsymbol{c}}\rangle\right)+\sum_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}\neq{\boldsymbol{c}}}\alpha_{{\boldsymbol{c}},{\boldsymbol{c}}^{\prime}}|{\boldsymbol{c}}-{\boldsymbol{c}}^{\prime}\rangle|Y_{{\boldsymbol{c}}^{\prime}}\rangle$ measure 1st register, if $|{\boldsymbol{0}}\rangle$: $\mapsto$ $\quad|\phi_{6}\rangle$ $=$ $\quad\dfrac{1}{\sqrt{|{\mathscr{C}}|}}\sum_{{\boldsymbol{c}}\in{\mathscr{C}}}|Y_{{\boldsymbol{c}}}\rangle=\dfrac{1}{Z_{0}}\sum_{{\boldsymbol{y}}\in{\mathscr{C}}^{\perp}}\hat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle\;\;\text{w.p. $\textrm{P}_{\textrm{PGM}}$}$ measuring the whole state: $\mapsto$ $\quad|{\boldsymbol{y}}\rangle\;\;\text{with ${\boldsymbol{y}}\in{\mathscr{C}}^{\perp}$.}$
@@
4.2 A tweak in the algorithm avoiding to measure the zero codeword
The idea is to tweak the algorithm in order to get as the last state before the final measurement the state $|U_{0}\rangle$ rather than $|\phi_{6}\rangle=\dfrac{1}{Z_{0}}\sum_{{\boldsymbol{y}}\in{\mathscr{C}}^{\perp}}\hat{f}({\boldsymbol{y}})|{\boldsymbol{y}}\rangle$, where
@@ 4.2 A tweak in the algorithm avoiding to measure the zero codeword
$|Y^{\prime}_{\boldsymbol{c}}\rangle=\frac{1}{\sqrt{q^{k}}}\left(|U_{0}\rangle+\sum_{{\boldsymbol{s}}\neq\mathbf{0}}\chi_{\boldsymbol{c}}(u_{\boldsymbol{s}})|\widetilde{W}_{\boldsymbol{s}}\rangle\right).$
These states are readily seen to be orthogonal. However projecting in this basis does not make the measurement complete. This can be circumvented by adding to the basis the $|{\boldsymbol{0}}\rangle$ state. We add an extra outcome $\boldsymbol{u}\in\mathbb{F}_{q}^{n}\backslash{\mathscr{C}}$, define ${\mathscr{S}}={\mathscr{C}}\cup\{\boldsymbol{u}\}$ and set $|Y^{\prime}_{\boldsymbol{u}}\rangle=|{\boldsymbol{0}}\rangle$. We therefore have a projective measurement $\{|Y^{\prime}_{{\boldsymbol{y}}}\rangle\}_{{\boldsymbol{y}}\in{\mathscr{S}}}$. Again, we have $\langle\widehat{\psi_{\boldsymbol{c}}}|{Y^{\prime}}_{\boldsymbol{c}}\rangle\geq\sqrt{\textrm{P}_{\textrm{PGM}}}-\frac{1}{\sqrt{q^{k}}}$ and independent of $\boldsymbol{c}$} so w.p. at least $\left(\sqrt{\textrm{P}_{\textrm{PGM}}}-\frac{1}{\sqrt{q^{k}}}\right)^{2}$, we get the state $|U_{0}\rangle$. Then, if we measure this state $|U_{0}\rangle$, we will get a dual codeword $\boldsymbol{y}$} with probability $q({\boldsymbol{y}})$ given by
@@
Let $q\geq 2$ be a prime power. We assume that we have an increasing sequence of positive integers $(n_{i})_{i\geq 1}$ and a family of functions $(f_{i})_{i\geq 1}$ with $f_{i}:\mathbb{F}_{q}^{n_{i}}\rightarrow\mathbb{C}$ with $\lvert\kern-1.0pt\lvert f_{i}\rvert\kern-1.0pt\rvert_{2}=1$ such that the sequence $(p_{i})_{i\geq 1}$ defined by $p_{i}\mathop{=}\limits^{\triangle}|\widehat{f_{i}}|^{2}$ admits a nice family of typical sets not containing 0. We let $H\mathop{=}\limits^{\triangle}\lim_{i\rightarrow\infty}\frac{H_{q}(p_{i})}{n_{i}}$. Let $R$ be a real number chosen in $(0,H)$. Consider a family $({\mathscr{C}}_{i})_{i\geq 1}$ of random linear codes where ${\mathscr{C}}_{i}$ is chosen with a generator matrix chosen uniformly at random in $\mathbb{F}_{q}^{\lfloor Rn_{i}\rfloor\times n_{i}}$. Then, for any $\epsilon>0$ the algorithm of the quantum reduction with PGM using the measurement basis $\{|Y^{\prime}_{{\boldsymbol{y}}}\rangle\}_{{\boldsymbol{y}}\in{\mathscr{S}}}$ described above, outputs codewords of $(\mathscr{C}_{i}^{\perp})^{*}$ in the typical set ${\mathcal{T}}_{\epsilon}^{(n_{i})}$} with probability $1-\operatorname{negl}(i)$ as $i$ tends to infinity.
@@
Let $g:{\mathbb{F}_{q}}\rightarrow\mathbb{C}$ with $\lvert\kern-1.0pt\lvert g\rvert\kern-1.0pt\rvert_{2}=1$ and $f:\mathbb{F}_{q}^{n}\rightarrow\mathbb{C}$ such that $f=g^{\otimes n}$ where $n$ is an arbitrary positive integer. Let $R$ be a real number in $(0,H_{q}(|\widehat{g}|^{2}))$. Choose a linear code $\mathscr{C}$ at random by picking its generator matrix uniformly at random in $\mathbb{F}_{q}^{\lfloor Rn\rfloor\times n}$. Then, for any $\epsilon>0$ the algorithm of the quantum reduction with PGM using the measurement basis $\{|Y^{\prime}_{{\boldsymbol{y}}}\rangle\}_{{\boldsymbol{y}}\in{\mathscr{S}}}$ described above, outputs codewords of $(\mathscr{C}^{\perp})^{*}$ in the typical set ${\mathcal{T}}_{\epsilon}^{(n)}$} with probability $1-\operatorname{negl}(n)$ as $n$ tends to infinity.